OpenStudy (vera_ewing):

1. Given the equation -4squrt(x-3)=12, solve for x and identify if it is an extraneous solution. x = 0, solution is not extraneous x = 0, solution is extraneous x = 12, solution is not extraneous x = 12, solution is extraneous

OpenStudy (vera_ewing):

@Directrix can you help with this one? the left side is negative but the right side is positive, so there can't be a solution...?

OpenStudy (vera_ewing):

@SolomonZelman please help?

OpenStudy (solomonzelman):

\(\large\color{black}{ 4\sqrt{x-3}=12 }\)

OpenStudy (vera_ewing):

-4

OpenStudy (solomonzelman):

\(\large\color{black}{ -4\sqrt{x-3}=12 }\) divide both sides by \(\large\color{black}{ -4 }\) .

OpenStudy (solomonzelman):

can you do that for me?

OpenStudy (solomonzelman):

delete this now!

OpenStudy (vera_ewing):

delete?

OpenStudy (solomonzelman):

not you

Directrix (directrix):

I got x = 12 but I have to run it back through the equation.

OpenStudy (solomonzelman):

\(\large\color{black}{ -4\sqrt{x-3}=12 }\) \(\large\color{black}{ \color{red}{(}-4\sqrt{x-3}\color{red}{)\div(-4)}=12\color{red}{)\div(-4)} }\)

OpenStudy (solomonzelman):

Directrix since when you are giving out answers like this?

OpenStudy (vera_ewing):

@Directrix but is it extraneous or not?

Directrix (directrix):

x = 12 is extraneous.

OpenStudy (vera_ewing):

ok thanks!

OpenStudy (amistre64):

when x=0 we get sqrt(-3), is this a valid operation when x=0?

OpenStudy (solomonzelman):

I am not reporting, lol...

Directrix (directrix):

It is extraneous and here is why.

Directrix (directrix):